### 12 Mar 2024 v1

# It’s All About Your Sketch: Democratising Sketch Control in Diffusion Models

Abstract
This paper unravels the potential of sketches for diffusion models, addressing the deceptive promise of direct sketch control in generative AI. We importantly democratise the process, enabling amateur sketches to generate precise images, living up to the commitment of “what you sketch is what you get”. A pilot study underscores the necessity, revealing that deformities in existing models stem from spatial-conditioning. To rectify this, we propose an abstraction-aware framework, utilising a sketch adapter, adaptive time-step sampling, and discriminative guidance from a pre-trained fine-grained sketch-based image retrieval model, working synergistically to reinforce fine-grained sketch-photo association. Our approach operates seamlessly during inference without the need for textual prompts; a simple, rough sketch akin to what you and I can create suffices! We welcome everyone to examine results presented in the paper and its supplementary. Contributions include democratising sketch control, introducing an abstraction-aware framework, and leveraging discriminative guidance, validated through extensive experiments.
1 Introduction
This paper is dedicated to unlocking the full potential of your sketches to control diffusion models [60, 23, 24]. Diffusion models  [60, 23, 24, 15] have made a significant impact, empowering individuals to unleash their visual creativity – consider prompts like “astronauts riding a horse on Mars" and other “creative” ones of your own! While prevailing in text-to-image generation [15, 63, 60], recent works [80, 90, 54] have started to question the expressive power of text as a conditioning modality. This shift has led to an exploration of sketches – a modality that offers a degree of fine-grained control that is unparalleled by text [12, 69], resulting in generated content of closer resemblance. The promise is “what you sketch is what you get”.
This promise is, however, deceptive. Current works (e.g., ControlNet [90], T2I-Adapter [54]) predominantly focus on curated edgemap-like sketches – you better sketch like a trained artist, otherwise “what you get” will literally be reflecting deformities captured in your (“half-decent”) sketch (LABEL:fig:teaser). The primary goal of this paper is to democratise sketch control in diffusion models, empowering real amateur sketches to generate photo-precise images, ensuring that “what you get” aligns with your intended sketch, regardless of how well you drew it! To achieve this, we draw insights from the sketch community [64, 87, 66, 37, 36] and introduce, for the first time, an awareness of sketch abstraction (as a result of varying drawing skills) into the generative process. This novel approach permits sketches of different abstraction levels to guide the generation process while maintaining output fidelity.
We conduct a pilot study to reaffirm the necessity of our research (Sec. 4). In which, we identify that the deformed output of existing sketch-conditional diffusion models stems from their spatial-conditioning approach – they directly translate sketch contours into the output photo domain, therefore producing deformed output. Conventional means of controlling the influence of spatial sketch-conditioning on the final output via weighing factors [54, 80] or sampling tricks [90], however, require careful tuning. Reducing output deformity by assigning less weight to the sketch-conditioning often makes the output more coherent with the textual description, thus reducing its fidelity to the guiding sketch; yet, assigning higher weight to the textual prompt introduces lexical ambiguity [70]. On the contrary, avoiding lexical ambiguity by assigning a higher weight to the guiding sketch almost always produces deformed and non-photorealistic outputs [90, 54, 80]. Last but not least, the sweet spot between the conditioning weights is different for different sketch instances (as seen in Fig. 1).
As such, our goal is to craft an effective sketch-conditioning strategy that not only operates without any textual prompts during inference but is also abstraction-aware. At the core of our work is a sketch adapter that transforms an input sketch into its equivalent textual embedding, directing the denoising process of the diffusion model via cross-attention. Through the use of a smart time-step sampling strategy, we ensure the adaptability of the denoising process to the abstraction level of the input sketch. Additionally, by capitalising on the pre-trained knowledge of an off-the-shelf [65] fine-grained sketch-based image retrieval (FG-SBIR) model, we incorporate discriminative guidance into our system for fine-grained sketch-photo association. Unlike widely used external classifier-guidance [15], our proposed discriminative guidance mechanism does not require any specifically trained classifier capable of classifying both noisy and real data. Lastly, even though our inference pipeline does not rely on textual prompts, we use synthetically generated textual prompts during training to learn the sketch adapter with the limited sketch-photo paired data.
Our contributions are: (i) we democratise sketch control, enabling real amateur sketches to generate accurate images, fulfilling the promise of “what you sketch is what you get”. (ii) we introduce an abstraction-aware framework that overcomes limitations of text prompts and spatial-conditioning. (iii) we leverage discriminative guidance through a pre-trained FG-SBIR model for fine-grained sketch-fidelity. Extensive experiments validate the effectiveness of our method in addressing existing limitations in this domain.
2 Related Works
Diffusion Model for Vision Tasks.   Diffusion models [24, 23, 73] have now become the gold-standard for different controllable image generation frameworks like DALL-E [56], Imagen [63], T2I-Adapter [54], ControlNet [90], etc. Besides image generation, several methods like Dreambooth [62], Imagic [31], Prompt-to-Prompt [21], SDEdit [51], SKED [53] extend it for realistic image editing. Beyond image generation and editing, diffusion model is also used in several downstream vision tasks like recognition [42], semantic [2] and panoptic [84] segmentation, image-to-image translation [78], medical imaging [14], image correspondence [77], image retrieval [38], etc.
Sketch for Visual Content Creation.  Following its success in sketch-based image retrieval (SBIR) [65, 10, 3], sketches are now being used in other downstream tasks like saliency detection [6], augmented reality [49, 50], medical image analysis [34], object detection [13], class-incremental learning [4], etc. Apart from the plethora of sketch-based $2D$ and $3D$ image generation and editing frameworks [35, 46, 59, 20, 90, 54, 80, 53, 81], sketches are also getting significant traction in other visual content creation tasks like animation generation [72] and inbetweening [71], creative garment design [45, 11], manga restoration [82], CAD modelling [43, 88], anime editing [27], etc.
Figure 1: Images generated by T2I-Adapter [54] for different sketch-guidance factors ($\omega\in[0,1]$). Determining the optimum $\omega$ to obtain an ideal balance (green-bordered) between photorealism and sketch-fidelity requires manual intervention and is sample-specific. A high value of $\omega$ works well for less deformed sketches, while the same for an abstract sketch produces deformed outputs and vice-versa.
Sketch-to-Image (S2I) Generation.  Prior GAN-based S2I models typically leverage either contextual loss [48], multi-stage generation [18], etc. or performs latent mapping [35, 59] on top of pre-trained GANs. Among diffusion-based frameworks, PITI [81] trains a dedicated encoder to map the guiding sketch to the pre-trained diffusion model’s latent manifold, SDEdit [51] sequentially adds noise to the guiding sketch and iteratively denoise it based on a text prompt, while SGDM [80] trains an MLP that maps the latent feature of the noisy images to the guiding sketches in order to force the intermediate noisy images to closely follow the guidance sketches. Among more recent multi-conditional (e.g., depth map, colour palate, key pose, etc.) frameworks, ControlNet [90] learns to control a frozen diffusion model by creating a trainable copy of its UNet encoders and connects it with the frozen model with zero-convolution [90], while T2I-Adapter [54] learns an encoder to extract features from the guidance signal (e.g., sketch) and conditions the generation process by adding the guidance features with the intermediate UNet features at each scale. While existing methods can generate photorealistic images from precise edgemaps, they struggle with abstract freehand sketches (see Fig. LABEL:fig:teaser). Furthermore, it is noteworthy that almost all of the diffusion-based S2I models [80, 81, 90, 54, 51] rely heavily on highly-engineered and detailed textual prompts.
3 Revisiting Diffusion Model (DM)
Overview.  Diffusion models comprises two complementary random processes viz.  “forward” and “reverse” [24] diffusion. Forward diffusion process iteratively adds Gaussian noise of varying magnitude to a clean training image $\mathbf{x}_{0}\in\mathbb{R}^{h\times w\times 3}$ for $t$ time-steps to yield a noisy image $\mathbf{x}_{t}\in\mathbb{R}^{h\times w\times 3}$ as:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+(\sqrt{1-\bar{\alpha}_{t}})\epsilon\vspace{-0.1cm}$ (1)
[…]
The unconditional denoising diffusion process could be made “conditional” by influencing the $\mathcal{F}_{\theta}$ with auxiliary conditioning signals ${d}$ (e.g., textual description [60, 57, 63], etc.). Thus, $\mathcal{F}_{\theta}(\mathbf{x}_{t},t,{d})$ could perform denoising on $\mathbf{x}_{t}$ while being guided by ${d}$ via cross-attention [60].
Latent Diffusion Model.  Unlike standard diffusion models [15, 24], Latent Diffusion Model [60] (a.k.a. Stable Diffusion–SD) performs denoising diffusion on the latent space for faster and more stable training [60]. SD first trains an autoencoder (consists of an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$ in series) to convert the input image $\mathbf{x}_{0}\in\mathbb{R}^{h\times w\times 3}$ to its latent representation $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}\times d}$. Later, SD trains a modified denoising UNet [61] $\epsilon_{\theta}(\cdot)$ to perform denoising directly on the latent space. The textual prompt ${d}$ upon passing through a CLIP textual encoder [55] $\mathbf{T}(\cdot)$ produces the corresponding token-sequence that influences the intermediate feature maps of the UNet via cross-attention [60]. SD trains with an $l_{2}$ loss as:
$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{d},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))||}_{2}^{2})$ (2)
During inference, SD discards $\mathcal{E}(\cdot)$, directly sampling a noisy latent $\mathbf{z}_{T}$ from a Gaussian distribution [60]. It then estimates noise from $\mathbf{z}_{T}$ iteratively for $T$ iterations via $\epsilon_{\theta}$ (conditioned on $d$) to obtain a clean latent $\hat{\mathbf{z}}_{0}$. The frozen decoder generates the final image as $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ [60].
4 What’s wrong with Sketch-to-Image DM
Recent controllable image generation methods like ControlNet [90], T2I-Adapter [54], etc. offer extreme photorealism, supporting different conditioning signals (e.g., depth map, label mask, edgemap, etc.) However, conditioning the same from sparse freehand sketches is often sub-optimal (LABEL:fig:teaser).
Sketch vs. Other Conditional Inputs. Sparse and binary freehand sketches while good for providing fine-grained spatial cues [89, 13, 6], often depict significant shape-deformity [22, 16, 64] and hold far less contextual information [78] than other pixel-perfect conditioning signals like depth maps, normal maps, or pixel-level segmentation masks. Hence, conditioning from freehand sketches is non-trivial and needs to be handled uniquely unlike the rest of the pixel-perfect conditioning signals.
Sketch vs. Text Conditioning: A Trade-off. Previous S2I diffusion models [54, 90, 80] exhibit two major challenges. Firstly, quality of generated outputs being highly dependent on precise and accurate textual prompts [90], inconsistencies or lack of suitable prompts can negatively impact (Fig. 2) the results [54, 90]. Secondly, ensuring a balance between the influence of sketch and text-conditioning on the final output requires manual intervention, which can be challenging. Adjusting the weighting of these factors often results in a trade-off between output’s coherence with the text and fidelity to the sketch [54]. In some cases, giving higher weight to text can lead to lexical ambiguity [70], while prioritising sketch tends to produce distorted and non-photorealistic results [54, 80]. Achieving photorealistic output from existing S2I DMs [54, 80] thus demands meticulous fine-tuning of these weights, where the optimal balance varies for different sketch instances as seen in Fig. 1.
[…]
Problems with Spatial-Conditioning for Sketches. We identify that the deformed and non-photorealistic (e.g., edge-bleeding in Fig. 1) outputs of existing sketch-conditional DMs [54, 90, 80] are primarily a consequence of their spatial-conditioning approach. T2I-Adapter [54] directly integrates the spatial features of the conditioning-sketch into the UNet encoder’s feature maps, while ControlNet [90] applies this to skip connections and middle blocks. SGDM [80], on the other hand, projects the latent features of noisy images to spatial edgemaps guiding the denoising process towards following the edgemaps. Additionally, these models are trained and tested with synthetically-generated [75, 7, 83] edgemaps/contours rather than real freehand sketches. Instead, we aim to devise an effective conditioning strategy for real freehand sketches while ensuring that the output faithfully captures an end-users’ semantic intent [35] without any deformities.
5 Proposed Methodology
Overview.  We aim to eliminate spatial sketch-conditioning by converting the input sketch into an equivalent fine-grained textual embedding, thereby preserving users’ semantic-intent without pixel-level spatial alignment. Consequently, our method would alleviate issues pertaining to spatial distortions (e.g., deformed shapes, edge-bleeding, etc.) while maintaining fine-grained fidelity to the input sketch. We introduce three salient designs (Fig. 3) – (i) fine-grained discriminative loss for maintaining the fine-grained sketch-photo correspondence (Sec. 5.2). (ii) guiding our training process with textual prompts (not used during inference), as a means of super-concept preservation (Sec. 5.3). Finally, (iii) unlike the uniform time-step ($t$) sampling of prior arts [90, 80], we introduce a sketch-abstraction-aware $t$-sampling (Sec. 5.4). For a highly abstract sketch, a higher probability is assigned to larger $t$ and vice-versa.
5.1 Sketch Adapter
Aiming to mitigate the evident disadvantages (Sec. 4) of direct spatial-conditioning approach of existing sketch-conditional diffusion models (e.g., ControlNet [90], T2I-Adapter [54], etc.), we take a parallel approach to “sketch-condition” the generation process via cross-attention. In that, instead of treating the input sketches spatially, we encode them as a sequence of feature vectors [41] as an equivalent fine-grained textual embedding. Direct spatial-conditioning enforces the model to remember the contextual information rather than understanding it [85]. This results in a direct translation of the strong sketch features (e.g., stroke boundaries) into the output photo. To overcome this, we aim to increase the hardness of the problem by compressing the spatial sketch input to a bottlenecked-representation via sketch adapter.
In particular, given a sketch $s$, we use a pre-trained CLIP [55] ViT-L/14 image encoder $\mathbf{V}(\cdot)$ to generate its patch-wise sketch embedding $\mathbf{s}=\mathbf{V}(s)\in\mathbb{R}^{257\times 1024}$. Our sketch adapter $\mathcal{A}(\cdot)$ consists of $1$-dimensional convolutional and vanilla attention [79] modules followed by FC layers. The convolutional and FC layers handle the dimension mismatch between text and sketch-embedding (i.e., $\mathbb{R}^{257\times 1024}$rightarrow $\mathbb{R}^{77\times 768}$), whereas the attention module tackles the large sketch-text domain gap. The patch-wise sketch embedding $\mathbf{s}$ upon passing through $\mathcal{A}(\cdot)$ generates the equivalent textual embedding as $\mathbf{\hat{s}}=\mathcal{A}(\mathbf{s})\in\mathbb{R}^{77\times 768}$. Now replacing the textual conditioning in Eq. 2 with our sketch adapter conditioning, the modified loss objective becomes:
$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{s},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||}_{2}^{2})\vspace{-0.1cm}$ (3)
Once trained, the sketch adapter efficiently converts an input sketch $s$ into its equivalent textual embedding $\hat{\mathbf{s}}$, which through cross-attention controls the denoising process of SD [60]. Nonetheless, conditioning solely via the proposed sketch adapter poses multiple challenges – (i) sparse freehand sketches and pixel-perfect photos depict a huge domain gap. The standard $l_{2}$ loss [60] of a text-to-image diffusion model is not enough to ensure a fine-grained matching between sketch and photo. (ii) training a robust sketch adapter from the limited available sketch-photo pairs is difficult. Consequently, during training, we aim to use pseudo texts as a learning signal to guide the training of our sketch adapter. Please note, our inference pipeline does not involve any textual prompts. (iii) the sketch adapter treats all sketch samples equally regardless of their abstraction levels. While this equal treatment might suffice for dense pixel-level conditioning, it might be inadequate for sparse sketches, as different sketches depicting different abstraction levels are not semantically-equal [5, 86].
Figure 3: Our overall training pipeline. (More in the text.)
5.2 Fine-Grained Discriminative Learning
To ensure a fine-grained matching between sparse freehand sketches and pixel-perfect photos, we utilise a pre-trained fine-grained (FG) SBIR model [65] $\mathcal{F}_{g}(\cdot)$. A photo sits close to its paired sketch in a pre-trained FG-SBIR model’s discriminative latent embedding space compared to other unpaired ones [65]. Previous attempts at guiding the diffusion process with external discriminative models include classifier-guidance [15] that require a pre-trained fixed-class classifier capable of classifying both noisy and real data [15] to guide the denoising procedure [15]. However, as our frozen FG-SBIR model is not trained on noisy data, it requires a clean image at each $t$, to perform in an off-the-shelf manner. Now, for each $t$, as the denoiser estimates that noise $\epsilon_{t}\approx\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))$, which was added to $\mathbf{z}_{0}$ to get $\mathbf{z}_{t}$ during forward diffusion, we can use Eq. 1 to recreate $\mathbf{z}_{0}$ from $\epsilon_{t}$. Specifically, we utilise Tweedie’s formula [33] to estimate [85, 1, 39] the clean latent image $\hat{\mathbf{z}}_{0}$ from the $t^{\text{th}}$-step noisy latent $\mathbf{z}_{t}$ in a single-step for efficient training as:
$\hat{\mathbf{z}}_{0}(\mathbf{z}_{t}):=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}~{}\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))}{\sqrt{\bar{\alpha}_{t}}}\vspace{-0.2cm}$ (4)
$\hat{\mathbf{z}}_{0}$ upon passing through SD’s [60] frozen VAE decoder $\mathcal{D}(\cdot)$ approximates the clean image $\hat{\mathbf{x}}_{0}$ (Sec. 3). To learn the sketch adapter $\mathcal{A}$, we use a discriminative SBIR loss that calculates cosine similarity $\delta(\cdot,\cdot)$ between $s$ and $\hat{\mathbf{x}}_{0}$ as:
$\mathcal{L}_{\text{SBIR}}=1-\delta\left({\mathcal{F}_{g}(s)\cdot\mathcal{F}_{g}(\hat{\mathbf{x}}_{0})}\right)\vspace{-0.1cm}$ (5)
5.3 Super-concept Preservation Loss
An inherent complementarity exists between sketch and text [12]. A textual caption of an image can correspond to multiple plausible photos in the embedding space. Adding a sketch with it however, narrows down the scope to a particular image [12, 69] (i.e., fine-grained). We posit that a textual description being less fine-grained than a sketch [12, 74, 85], acts as a super-concept of the corresponding sketch. Although we do not use any textual prompt during inference, we aim to use them during training of our sketch adapter. Text-to-image diffusion models being trained on a large corpus of text-image pairs [60], implicitly hold superior text-to-image generation capability (although not fine-grained [17]). We thus aim to use this super-concept knowledge from textual descriptions to distil the large-scale text-to-image knowledge of a pre-trained SD to train our sketch-adapter with limited sketch-photo paired data.
As our sketch-photo ($s,p$) dataset [68] lacks paired textual captions, we use a pre-trained state-of-the-art image captioner [44] to synthetically generate caption $d$ for every ground truth photo $p$. Then, at each $t$, the noise predicted through text-conditioning $(\mathbf{T}(d))$ acts as a reference to calculate a regularisation loss to learn the sketch adapter $\mathcal{A}$ as:
$\mathcal{L}_{\mathrm{reg}}=||\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||_{2}^{2}\vspace{-0.1cm}$ (6)
5.4 Abstraction-aware Importance Sampling
[…]
$\mathcal{S}_{\omega}(t)=\frac{1}{T}\left(1-\omega~{}\mathrm{cos}\frac{\pi t}{T}\right)\vspace{-0.1cm}$ (7)
where, $\mathcal{S}_{\omega}(\cdot)$ is our abstraction-aware $t$-sampling function, where increasing or decreasing $\omega\in(0,1]$, controls the skewness of this sampling probability density function. Pushing $\omega$ towards $1$ increases the probability of sampling a larger $t$ value. We aim to make this skewness-controlling $\omega$ value sketch-abstraction specific (Fig. 4).
Figure 4: Abstraction-aware $t$-sampling function for different $\omega$.
Now the question remains as to how we can quantify the abstraction level of a freehand sketch. Taking inspiration from [87], we design a CLIP [55]-based (a generic classifier) sketch classifier with a MagFace [52]-based loss where the $l_{2}$-norm of a sketch feature $\mathbf{a}\in[0,1]$, denotes how closely it sits from its respective class-centre. While $\mathbf{a}$rightarrow $1$ represents edgemap-like less abstract sketches, $\mathbf{a}$rightarrow 00 denotes highly-abstract and deformed ones. We posit that edgemaps being less deformed (i.e., easier to classify), will implicitly stay close to their respective class centres in the latent space. Whereas, freehand sketches being highly abstract and deformed (i.e., harder to classify), will be placed away from their corresponding class centres. We thus train the sketch classifier with sketches and synthesised [9] edgemaps of the associated photos from Sketchy [68], using our classification loss:
$\footnotesize\mathcal{L}_{\mathrm{abs}}=-\mathrm{log}\frac{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}}{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}+\sum_{j\neq y_{i}}e^{s~{}\mathrm{cos}~{}\theta_{j}}}+\lambda_{g}g(\mathbf{s}_{i})$ (8)
where $s$ is a global scalar value, $\theta_{y_{i}}$ is the cosine similarity between extracted global visual feature (from CLIP [55] visual encoder) of the $i^{\text{th}}$ sketch sample $\mathbf{s}_{i}=\mathbf{V}(s_{i})\in\mathbb{R}^{d}$ with $l_{2}$-normalisation, and $j^{\text{th}}$ class centre $w_{j}\in\mathbb{R}^{d}$ computed from ground truth class labels by CLIP [55] text encoder. $m(\mathbf{s}_{i})$ is the magnitude-aware margin parameter $m(\mathbf{s}_{i})=\frac{(u_{m}-l_{m})}{(u_{a}-l_{a})l_{a}+l_{m}}$, where $l_{m}$, $u_{m}$ denotes the lower and upper bounds of the margin, and $l_{a}$, $u_{a}$ denotes that of the feature magnitude. $g(\mathbf{s}_{i})$ is a hyper-parameter ($\lambda_{g}$)-controlled regularisation term (see [52] for more details). With the trained classifier, given a sketch $s$, the scalar abstraction score $\mathbf{a}\in[0,1]$ is given by the $l_{2}$-norm of the extracted sketch feature $\mathbf{V}(s)$. To keep parity with $\omega$, we complement $\mathbf{a}$ to get the sketch instance-specific $\omega\leftarrow(1-\mathbf{a})$, followed by empirically clipping $\omega$ in the range $[0.2,0.8]$.
In summary, we train the sketch adapter $\mathcal{A}(\cdot)$ using sketch-abstraction-aware $t$-sampling with a total loss of $\mathcal{L}_{\mathrm{total}}$=$\lambda_{1}\mathcal{L}_{\mathrm{SD}}$+$\lambda_{2}\mathcal{L}_{\mathrm{SBIR}}$+$\lambda_{3}\mathcal{L}_{\mathrm{reg}}$. During inference, we compute the abstraction score of the input sketch, taking $l_{2}$-norm of classifier feature. Based on the abstraction level, we perform $t$-sampling. The input sketch passing through $\mathcal{A}$ controls the diffusion procedure and generates the output.
6 Experiments
Dataset and Implementation Details.   We train and evaluate our model on the Sketchy dataset [68] containing $12,500$ images from $125$ categories with at least $5$ sketches per image with fine-grained association. For training and evaluation, we split this dataset in $90$:$10$. We use Stable Diffusion v1.5 [60] in all experiments with a CLIP [55] embedding dimension $d=768$. The sketch adapter is trained with a learning rate of $10^{-4}$, keeping the SD model, FG-SBIR backbone, and CLIP encoders frozen. We train our model for $50$ epochs using AdamW [47] optimiser with $0.09$ weight decay, and batch size of $8$. Values of $\lambda_{1,2,3}$ are set to $1$, $0.5$, and $0.1$, empirically.
Evaluation Metrics.  Following [54, 90, 35], we quantitatively evaluate the generation quality and sketch-fidelity with four metrics – Frechèt Inception Distance-InceptionV3 (FID-I) [30] and CLIP (FID-C) [40] calculates the similarity between generated and real images using pre-trained InceptionV3 [76] and CLIP [55] ViT-B/32 models respectively. Lower values of FID-I and FID-C depict better generation quality. We measure the output image’s fidelity to the input sketch using Fine-Grained Metric (FGM) [35] which computes the cosine similarity between them via a pre-trained FG-SBIR model [65], where a higher value denotes better fine-grained correspondence. Additionally, we also perform a human study to collect Mean Opinion Score (MOS) [28]. Here, we asked $25$ non-artist users to draw $40$ sketches each, and rate the generated photos on a discrete scale (interval=$0.5$) of $[1,5]$ (worst to best) based on output photorealism and sketch-fidelity. For each method, we compute the final MOS by averaging all its $1000$ MOS values.
[…]
6.1 Performance Analysis & Discussion
Result Analysis.  Among GAN-based methods, pix2pix [29] and CycleGAN [91] depict visible deformities (Fig. 5) mostly due to their weaker [15] GAN-based generator, compared to an internet-scale pre-trained SD model [60]. Among diffusion-based SOTAs, although SGDM [80] generates plausible colour schemes and styles, outputs exhibit substantial deformations (LABEL:fig:teaser). A similar observation can be made for PITI [81], where generated images look non-photorealistic with pronounced edge-adherence (Fig. 5). Whereas, edge-bleeding (Fig. 5) is quite frequent for T2I-Adapter [54]. ControlNet [90] surpasses PITI [81], SGDM [80], and T2I-Adapter [54] in terms of photorealism but mostly follows the input sketch boundaries (Fig. 5). Contrarily, images generated by our method are more photorealistic with fewer deformities, capturing semantic-intent without transmitting edge boundaries in the output. Quantitative results presented in Tab. 1 show B-Caption to surpass B-Classification (by $0.11$ FGM) thanks to the comparatively higher [44] generalisation potential of the captioning model [44] than the generic sketch classifier [32]. Nonetheless, our method exceeds these baselines both in terms of generation quality and sketch-fidelity with an FID-C of $16.20$ and FGM of $0.81$. Due to its superior conditioning strategy, ControlNet [90] achieves the lowest FID-I among all prior SOTAs (Tab. 1). Although less pronounced in terms of FID-I/FID-C, our method offers the highest fine-grained sketch-fidelity with $23.45\%$ FGM improvement of over ControlNet [90]. Finally, thanks to the photorealistic generation quality and fine-grained sketch correspondence, our method surpasses competitors in terms of MOS value from user-study with an average $1.36\pm 0.2$ point improvement. Notably, unlike ours, image generation via diffusion-based competitors needs textual prompts, the absence of which results in much worse output quality (Fig. 2).
Table 1: Benchmarks on the Sketchy [68] dataset.
Methods FID-I $\downarrow$ FID-C $\downarrow$ FGM $\uparrow$ MOS $\uparrow$
$\mu\pm\sigma$
ControlNet [90] 26.68 21.22 0.62 3.68$\pm$0.2
T2I-Adapter [54] 26.94 18.92 0.56 3.11$\pm$0.6
PITI [81] 84.71 25.85 0.23 2.64$\pm$0.3
B-Classification 28.93 19.01 0.36 3.13$\pm$0.2
B-Captioning 28.31 18.81 0.47 3.21$\pm$0.4
Proposed 25.07 16.20 0.81 4.52$\pm$0.1
Generalisation Potential.  As our method alleviates the direct spatial influence of input sketches in the denoising process, it enables generalisation across multiple dimensions. Fig. 6 shows that our sketch-adapter trained on Sketchy, generalises well on random sketch samples from TU-Berlin [16] and QuickDraw [19] datasets, on synthetically generated [7] edgemaps, and to different stroke-styles. Furthermore, as our sketch adapter does not distort the original text-to-image pre-training of the frozen SD model, the same adapter could be used to perform sketch-conditional generation from other versions of the SD model (Fig. 7).
Figure 6: Examples showing generalisation potential across different datasets (left) and stroke-styles (right).
Figure 7: Illustration of cross-model generalisation. Our method trained with SD v1.5 [60], performs well on other unseen SD variants (e.g., v1.4) without further fine-tuning.
Robustness and Sensitivity.  Amateur freehand sketching often introduces irrelevant and noisy strokes [5]. We thus demonstrate our model’s resilience to such strokes by progressively adding them during inference, and assessing its performance. On the other hand, to judge our model’s stability against partially-complete sketches, we render input sketches at $\{25,50,75,100\}\%$ prior to generation. As our method is devoid of direct spatial-conditioning, outputs remain relatively stable (Fig. 8) even for spatially distorted sketches (e.g., noisy or partially-complete).
Figure 8: Examples depicting the effect of adding noisy strokes (left) and generation from partially-completed sketches (right).
Fine-grained Semantic Editing.  Harnessing the large-scale pre-training of the frozen SD model [60], our method enables fine-grained semantic editing. Here, fixing the generation seed, and performing local semantic edits in the sketch-domain produces seamless edited images (Fig. 9).
Figure 9: Our method seamlessly transfers local semantic edits on input sketches into output photos. (Best view when zoomed in.)
6.2 Ablation on Design
[i] Importance of Sketch Adapter. Our sketch adapter (Sec. 5.1) converts an input sketch to its corresponding textual equivalent embedding. To judge its efficacy, we replace it with simple convolutional and FC-layers converting the $\mathbb{R}^{257\times 1024}$ sketch embedding to equivalent $\mathbb{R}^{77\times 768}$ textual embedding. Although less pronounced in FID scores, the FGM score plummets substantially ($49.38\%$) in case of w/o Sketch adapter (Tab. 2), indicating the significance of the proposed adapter in maintaining high sketch-fidelity.
[ii] Why Discriminative Learning? Fine-grained discriminative loss (Eq. 5) helps the conditioning process by distilling knowledge learned inside a pre-trained FG-SBIR model. As seen in Tab. 2, a noticeable FGM drop ($44.44\%$) for w/o Discriminative learning indicates that fine-grained sketch-conditioning is incomplete without explicit discriminative learning via $\mathcal{L}_{\mathrm{SBIR}}$.
[iii] Does Abstraction-aware Importance Sampling help? Unlike existing sketch-conditional DMs, we take freehand sketch abstraction into account via abstraction-aware $t$-sampling. Omitting it results (Tab. 2) in a sharp increase in FID-I scores ($26.64\%$). We hypothesise that in absence of the proposed adaptive $t$-sampling, the system treats all sketches equally, regardless of their abstraction level, resulting in sub-optimal performance.
[iv] Impact of Super-concept Preservation. Although our inference procedure does not use any textual prompt, we employ them during our training process to facilitate the preservation of super-concepts. Eliminating this again destabilises the system causing an additional $15.06\%$ and $17.28\%$ decline in FID-C and FGM scores (Tab. 2). This justifies our incorporation of synthetic text prompts during training, as it aligns well with the original text-to-image generation objective of the pre-trained SD model [60]. Visual ablation results are presented in Fig. 10.
Figure 10: Visual ablation of different design components.
Table 2: Ablation on design.
Methods FID-I $\downarrow$ FID-C $\downarrow$ FGM $\uparrow$
w/o Sketch adapter 29.23 20.34 0.41
w/o Discriminative learning 29.14 19.97 0.45
w/o Super-concept preservation 27.21 18.64 0.67
w/o Abs.-aware $t$-sampling 31.75 23.17 0.55
Ours (SD v1.4) 26.12 17.09 0.77
Ours-full 25.07 16.20 0.81
6.3 Failure Cases & Future Works
Despite showcasing superior generation quality without significant deformations, our method has a few limitations. For Instance, it sometimes struggles to determine the correct class of the input due to categorical-ambiguity, especially when two different objects look very similar shape-wise (Fig. 11) in their abstract and deformed sketch forms (e.g., apple vs. pear, guitar vs. violin). In future, we aim to extend our method with the flexibility to include additional class labels. The sketch+label composed-conditioning [67] might mitigate the categorical-ambiguity of confusing classes.
Figure 11: Failure cases where sketches from certain classes (e.g., zebra) produce images from other similar-looking classes (e.g., horse) or vice-versa. Please note that we do not use text prompts.
7 Conclusion
Our work takes a significant step towards democratising sketch control in diffusion models. We exposed the limitations of current approaches, showcasing the deceptive promise of sketch-based generative AI. By introducing an abstraction-aware framework, featuring a sketch adapter, adaptive time-step sampling, and discriminative guidance, we empower amateur sketches to yield precise, high-fidelity images without the need for textual prompts during inference. We welcome the community to scrutinise our results. Please refer to the demo video for a detailed real-time comparison with state-of-the-arts.
Supplementary material for
It’s All About Your Sketch: Democratising Sketch Control in Diffusion Models
Subhadeep Koley1,2    Ayan Kumar Bhunia1    Deeptanshu Sekhri1    Aneeshan Sain1,2
Pinaki Nath Chowdhury1,2    Tao Xiang1,2    Yi-Zhe Song1,2
1SketchX, CVSSP, University of Surrey, United Kingdom.
2iFlyTek-Surrey Joint Research Centre on Artificial Intelligence.
{s.koley, a.bhunia, d.sekhri, a.sain, p.chowdhury, t.xiang, y.song}@surrey.ac.uk
A. Additional Qualitative Results
Fig. 12 delineates qualitative comparison of our method with pix2pix [29], CycleGAN [91], ControlNet [90], T2I-Adapter [54], and PITI [81]. Whereas, Fig. 13-18 shows additional results generated by our method.
[…]
Figure 13: Images generated by our method.
Figure 14: Images generated by our method.
Figure 15: Images generated by our method.
Figure 16: Images generated by our method.
Figure 17: Images generated by our method.
Figure 18: Images generated by our method.
B. Results on Out-of-distribution Sketches
Keeping the pre-trained diffusion model frozen, we fully leverage its generalisation potential. We posit that our design enables out-of-distribution generalisation. Fig. 19 shows a few sketches, absent in Sketchy [68].
Figure 19: Results on out-of-distribution sketches.

### 20 Mar 2024 v2

@@
 2 Related Works
-Diffusion Model for Vision Tasks.
+Diffusion Models for Vision Tasks.
 Diffusion models [25, 24, 74] have now become the gold-standard for different controllable image generation frameworks like DALL-E [57], Imagen [64], T2I-Adapter [55], ControlNet [90], etc.
@@
 Following its success in sketch-based image retrieval (SBIR) [66, 11, 3], sketches are now being used in other downstream tasks like saliency detection [6], augmented reality [50, 51], medical image analysis [35], object detection [14], class-incremental learning [4], etc.
-Apart from the plethora of sketch-based $2D$ and $3D$ image generation and editing frameworks [35, 46, 59, 20, 90, 54, 80, 53, 81], sketches are also getting significant traction in other visual content creation tasks like animation generation [72] and inbetweening [71], creative garment design [45, 11], manga restoration [82], CAD modelling [43, 88], anime editing [27], etc.
+Apart from the plethora of sketch-based $2D$ and $3D$ image generation and editing frameworks [36, 47, 60, 21, 90, 55, 81, 54, 82], sketches are also getting significant traction in other visual content creation tasks like animation generation [73] and inbetweening [72], garment design [46, 12], caricature generation [10], CAD modelling [44, 88], anime editing [28], etc.
 Figure 1: Images generated by T2I-Adapter [55] for different sketch-guidance factors ($\omega\in[0,1]$).
@@
 SD trains with an $l_{2}$ loss as:
-$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{d},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))||}_{2}^{2})$ (2)
+$\mathcal{L}_{\text{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{d},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))||}_{2}^{2})$ (2)
 During inference, SD discards $\mathcal{E}(\cdot)$, directly sampling a noisy latent $\mathbf{z}_{T}$ from a Gaussian distribution [61].
 It then estimates noise from $\mathbf{z}_{T}$ iteratively for $T$ iterations via $\epsilon_{\theta}$ (conditioned on $d$) to obtain a clean latent $\hat{\mathbf{z}}_{0}$.
-The frozen decoder generates the final image as $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ [60].
+The frozen decoder generates the final image as: $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ [61].
 4 What’s wrong with Sketch-to-Image DM
@@
 2 with our sketch adapter conditioning, the modified loss objective becomes:
-$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{s},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||}_{2}^{2})\vspace{-0.1cm}$ (3)
+$\mathcal{L}_{\text{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{s},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||}_{2}^{2})\vspace{-0.1cm}$ (3)
 Once trained, the sketch adapter efficiently converts an input sketch $s$ into its equivalent textual embedding $\hat{\mathbf{s}}$, which through cross-attention controls the denoising process of SD [61].
@@
 Then, at each $t$, the noise predicted through text-conditioning $(\mathbf{T}(d))$ acts as a reference to calculate a regularisation loss to learn the sketch adapter $\mathcal{A}$ as:
-$\mathcal{L}_{\mathrm{reg}}=||\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||_{2}^{2}\vspace{-0.1cm}$ (6)
+$\mathcal{L}_{\text{reg}}=||\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||_{2}^{2}\vspace{-0.1cm}$ (6)
 5.4 Abstraction-aware Importance Sampling
@@
 where, $\mathcal{S}_{\omega}(\cdot)$ is our abstraction-aware $t$-sampling function, where increasing or decreasing $\omega\in(0,1]$, controls the skewness of this sampling probability density function.
-Pushing $\omega$ towards $1$ increases the probability of sampling a larger $t$ value.
-We aim to make this skewness-controlling $\omega$ value sketch-abstraction specific (Fig.
+Pushing $\omega$ towards $1$ increases the probability of sampling a larger $t$ value (Fig.
 4).
+We aim to make this skewness-controlling $\omega$ value sketch-abstraction specific.
 Figure 4: Abstraction-aware $t$-sampling function for different $\omega$.
@@
 We thus train the sketch classifier with sketches and synthesised [9] edgemaps of the associated photos from Sketchy [69], using our classification loss:
-$\footnotesize\mathcal{L}_{\mathrm{abs}}=-\mathrm{log}\frac{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}}{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}+\sum_{j\neq y_{i}}e^{s~{}\mathrm{cos}~{}\theta_{j}}}+\lambda_{g}g(\mathbf{s}_{i})$ (8)
+$\footnotesize\mathcal{L}_{\text{abs}}=-\mathrm{log}\frac{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}}{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}+\sum_{j\neq y_{i}}e^{s~{}\mathrm{cos}~{}\theta_{j}}}+\lambda_{g}g(\mathbf{s}_{i})$ (8)
 where $s$ is a global scalar value, $\theta_{y_{i}}$ is the cosine similarity between extracted global visual feature (from CLIP [56] visual encoder) of the $i^{\text{th}}$ sketch sample $\mathbf{s}_{i}=\mathbf{V}(s_{i})\in\mathbb{R}^{d}$ with $l_{2}$-normalisation, and $j^{\text{th}}$ class centre $w_{j}\in\mathbb{R}^{d}$ computed from ground truth class labels by CLIP [56] text encoder.
@@
 To keep parity with $\omega$, we complement $\mathbf{a}$ to get the sketch instance-specific $\omega\leftarrow(1-\mathbf{a})$, followed by empirically clipping $\omega$ in the range $[0.2,0.8]$.
-In summary, we train the sketch adapter $\mathcal{A}(\cdot)$ using sketch-abstraction-aware $t$-sampling with a total loss of $\mathcal{L}_{\mathrm{total}}$=$\lambda_{1}\mathcal{L}_{\mathrm{SD}}$+$\lambda_{2}\mathcal{L}_{\mathrm{SBIR}}$+$\lambda_{3}\mathcal{L}_{\mathrm{reg}}$.
+In summary, we train the sketch adapter $\mathcal{A}(\cdot)$ using sketch-abstraction-aware $t$-sampling with a total loss of $\mathcal{L}_{\text{total}}$=$\lambda_{1}\mathcal{L}_{\text{SD}}$+$\lambda_{2}\mathcal{L}_{\text{SBIR}}$+$\lambda_{3}\mathcal{L}_{\text{reg}}$.
 During inference, we compute the abstraction score of the input sketch, taking $l_{2}$-norm of classifier feature.
@@
 As seen in Tab.
-2, a noticeable FGM drop ($44.44\%$) for w/o Discriminative learning indicates that fine-grained sketch-conditioning is incomplete without explicit discriminative learning via $\mathcal{L}_{\mathrm{SBIR}}$.
+2, a noticeable FGM drop ($44.44\%$) for w/o Discriminative learning indicates that fine-grained sketch-conditioning is incomplete without explicit discriminative learning via $\mathcal{L}_{\text{SBIR}}$.
 [iii] Does Abstraction-aware Importance Sampling help?
@@
 It’s All About Your Sketch: Democratising Sketch Control in Diffusion Models
-Subhadeep Koley1,2    Ayan Kumar Bhunia1    Deeptanshu Sekhri1    Aneeshan Sain1,2
-Pinaki Nath Chowdhury1,2    Tao Xiang1,2    Yi-Zhe Song1,2
+Subhadeep Koley1,2    Ayan Kumar Bhunia1    Deeptanshu Sekhri1    Aneeshan Sain1
+Pinaki Nath Chowdhury1    Tao Xiang1,2    Yi-Zhe Song1,2
 1SketchX, CVSSP, University of Surrey, United Kingdom.